Shawn and his bike have a total mass of 44.2kg.Shawn rides his bike 0.73km in 13.2 min at a constant velocity. The acceleration of gravity is 9.8 m/s^2. What is Shawn's kinetic ? Answer in units of J
1
Expert's answer
2018-11-19T16:01:10-0500

K=1/2 mv^2

K=1/2 (44.2) (730/(13.2∙60))^2=18.8 J.